# Environment-Driven Online LiDAR-Camera Extrinsic Calibration

### v1

@@ Abstract
LiDAR-camera extrinsic calibration (LCEC) is the core for data fusion in computer vision. Existing methods typically rely on customized calibration targets or fixed scene types, lacking the flexibility to handle variations in sensor data and environmental contexts. This paper introduces EdO-LCEC, the first environment-driven, online calibration approach that achieves human-like adaptability. Inspired by the human perceptual system, EdO-LCEC incorporates a generalizable scene discriminator to actively interpret environmental conditions, creating multiple virtual cameras that capture detailed spatial and textural information. To overcome cross-modal feature matching challenges between LiDAR and camera, we propose dual-path correspondence matching (DPCM), which leverages both structural and textural consistency to achieve reliable 3D-2D correspondences. Our approach formulates the calibration process as a spatial-temporal joint optimization problem, utilizing global constraints from multiple views and scenes to improve accuracy, particularly in sparse or partially overlapping sensor views. Extensive experiments on real-world datasets demonstrate that EdO-LCEC achieves state-of-the-art performance, providing reliable and precise calibration across diverse, challenging environments.
@@ I Introduction
We have long envisioned robots with human-like intelligence, enabling them to understand, adapt to, and positively impact the world [1, 2, 3, 4]. This dream is becoming increasingly attainable with the advent of LiDAR-camera data fusion systems. LiDARs provide accurate spatial information, while cameras capture rich textural details [5]. The complementary strengths of these two modalities, when fused, provide robots with powerful environmental perception capabilities [6, 7]. LiDAR-camera extrinsic calibration (LCEC), which estimates the rigid transformation between the two sensors, is a core and foundational process for effective data fusion.
@@
Current LiDAR-camera extrinsic calibration methods are primarily categorized as either target-based or target-free. Target-based approaches [8, 9, 10, 11, 12, 13] have long been the preferred choice in this field. They are typically offline, relying on customized calibration targets (typically checkerboards). However, they often demonstrate poor adaptability to real-world environments. This is largely because extrinsic parameters may change significantly due to moderate shocks or during extended operations in environments with vibrations. Online, target-free approaches aim at overcoming this problem by extracting informative visual features directly from the environment. Previous works [14, 15] estimate the extrinsic parameters by matching the cross-modal edges between LiDAR projections and camera images. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed edge features. Recent advances in deep learning techniques have spurred extensive explorations, such as [16] and [5], to leverage semantic information to aid cross-modal feature matching. Although these approaches have shown compelling performance in specific scenarios, such as parking lots, they predominantly rely on curated, pre-defined objects, e.g., vehicles [17], lanes [16], poles [18], and stop signs [19]. On the other hand, several end-to-end deep learning networks [20, 21, 22, 23, 24, 25] have been developed to find a more direct solution for LCEC. While these methods have demonstrated effectiveness on large-scale datasets like KITTI [26], they suffer from a high dependency on the training setup and are thus less generalizable. Overall, existing methods passively rely on predefined techniques to identify certain preset scene features. They lack the flexibility needed to adapt and respond effectively to environmental changes. In contrast, human actions are environment-driven. People intuitively assess their surroundings, swiftly respond to changes, make adaptive decisions, and improve through interaction and feedback. Elevating online calibration to this level of human-like adaptability presents a compelling research direction.
@@
In this work, we take one step forward in the field of online LCEC by designing an environment-driven framework EdO-LCEC. As illustrated in Fig. Environment-Driven Online LiDAR-Camera Extrinsic Calibration, our approach treats sensors’ operational environment as a spatial-temporal flow. By selectively combining information from multiple scenes across different times, this approach could achieve high-precision calibration dynamically. A prerequisite for the success of EdO-LCEC is the design of a generalizable scene discriminator. The scene discriminator employs large vision models to conduct depth estimation and image segmentation, achieving high-quality environmental perception on semantic and spatial dimensions. In detail, it calculates the feature density of the calibration scene and uses this information to guide the generation of multiple virtual cameras for projecting LiDAR intensities and depth. This improved LiDAR point cloud projecting strategy increases the available environmental features, and thus overcomes the previous reliance of algorithms [14, 27] on uniformly distributed geometric or textural features. At each scene, we perform dual-path correspondence matching (DPCM). Drawing inspiration from the human visual system [28], DPCM divides correspondence matching into spatial and textural pathways to fully leverage both geometric and semantic information. Each pathway constructs a cost matrix based on structural and textural consistency, guided by accurate semantic priors, to yield reliable 3D-2D correspondences. Finally, the correspondences obtained from multiple views and scenes are used as inputs for our proposed spatial-temporal relative pose optimization, which derives and refines the extrinsic matrix between LiDAR and camera. Through extensive experiments conducted on three real-world datasets, EdO-LCEC demonstrates superior robustness and accuracy compared to other SoTA online, target-free approaches.
@@
EdO-LCEC, the first environment-driven, online LCEC framework that achieves high calibration accuracy and human-like adaptability.
Generalizable scene discriminator, which can automatically perceive the calibration scene and capture potential spatial, textural and semantic features using SoTA LVMs.
@@
Spatial-temporal relative pose optimization, enabling high-quality extrinsic estimation through multi-view and multi-scene optimization.
@@ II Related Work
Target-based methods achieve high accuracy using customized calibration targets (typically checkerboards). However, they require offline execution and are significantly limited in dynamic or unstructured environments where such targets are unavailable [29, 30]. Recent studies have shifted to online, target-free methods to overcome these limitations. Pioneering works [31, 32, 14, 15, 33, 34] estimate the relative pose between the two sensors by aligning the cross-modal edges or mutual information (MI) extracted from LiDAR projections and camera RGB images. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed edges and rich texture, which largely compromise calibration robustness. To circumvent the challenges associated with cross-modal feature matching, several studies [35, 36, 37, 29] have explored motion-based methods. These approaches match sensor motion trajectories from visual and LiDAR odometry to derive extrinsic parameters through optimization. While they effectively accommodate heterogeneous sensors without requiring overlap, they demand precise synchronized LiDAR point clouds and camera images to accurately estimate per-sensor motion, which limits their applicability.
Advances in deep learning techniques have driven significant exploration into enhancing traditional target-free algorithms. Some studies [17, 16, 18, 19, 30, 38, 27, 5] explore attaching deep learning modules to their calibration framework as useful tools to enhance calibration efficiency. For instance, [16] accomplishes LiDAR and camera registration by aligning road lanes and poles detected by semantic segmentation. Similarly, [19] employs stop signs as calibration primitives and refines results over time using a Kalman filter. A recent study [27] introduced Direct Visual LiDAR Calibration (DVL), a novel point-based method that utilizes SuperGlue [39] to establish direct 3D-2D correspondences between LiDAR and camera data. On the other hand, several learning-based algorithms [20, 21, 22, 23, 24, 25] try to covert calibration process into more direct approaches by exploiting end-to-end deep learning networks. Although these methods have demonstrated effectiveness on public datasets like KITTI [26], which primarily focuses on urban driving scenarios, their performance has not been extensively validated on other types of real-world datasets that contain more challenging scenes. Furthermore, their dependence on pre-defined sensor configurations (both LiDAR and camera) poses implementation challenges. In comparison, our approach actively perceives the environment and achieves human-like adaptability. This environment-driven strategy enables robots to rapidly obtain high-precision extrinsic parameters anytime and anywhere.
@@ III Methodology
We first utilize a scene discriminator to understand the environment through image segmentation and depth estimation, generating virtual cameras that project LiDAR intensity and depth from multiple viewpoints (Sect. III-A).
The image segmentation outputs are processed along two pathways (spatial and textural), then input into a dual-path correspondence matching module to establish reliable 3D-2D correspondences (Sect. III-B).
The obtained correspondences are used as inputs for our proposed spatial-temporal relative pose optimization, which derives and refines the extrinsic matrix (Sect. III-C).
@@ III-A Generalizable Scene Discriminator
Our environment-driven approach first employs a generalizable scene discriminator to interpret the surroundings by generating virtual cameras to project LiDAR point cloud intensities and depth. This discriminator configures both an intensity and a depth virtual camera from the LiDAR’s perspective. This setup yields a LiDAR intensity projection (LIP) image ( and represent the image height and width) and a LiDAR depth projection (LDP) image . To align with the LDP image, the input camera RGB image is processed using Depth Anything V2 [40] to obtain estimated depth images 111In this article, the symbols in the superscript denote the type of target camera ( denotes virtual camera and indicates real camera), and the subscript denotes the source of the image ( is depth and is intensity).. To take advantage of semantic information, we utilize MobileSAM [41] as the image segmentation backbone. The series of detected masks in an image is defined as . The corner points along the contours of masks detected are represented by , where is the total corner points number in the -th mask. An instance (bounding box), utilized to precisely fit around each mask, is centrally positioned at and has a dimension of pixels. We define the corresponding depth of a corner point as , and as the matrix that stores texture values in its -neighborhood. To fully leverage the structured information within the masks, we construct a convex polygon from the corner points. As depicted in Fig. 1, the neighboring vertices of a corner point is defined as .
@@
where denotes the textural density and represents the structural density. The occlusion challenges caused by the different perspectives of LiDAR and the camera [14, 5], combined with limited feature availability in sparse point clouds, mean that a single pair of virtual cameras is insufficient for a comprehensive view of the calibration scene. Let represent the event of capturing enough effective features, with probability . If we have intensity virtual cameras and depth virtual cameras, the probability of capturing enough effective features is . In theory, as , the probability , leading to . Increasing the number of virtual cameras raises the likelihood of capturing more potential correspondences, thus enhancing calibration accuracy. However, applying an infinite number of virtual cameras during calibration is impractical. Considering the trade-off between calibration accuracy and compute resources, we set the number of multiple virtual cameras to satisfy the balance of feature density:
In practical applications, we set multiple virtual cameras inside a sphere originating from the LiDAR perspective center. Since the feature density is similar if the perspective viewpoints of virtual cameras are close to the initial position, we can assume that and . So and can be obtained as follows:
@@ III-B Dual-Path Correspondence Matching
Given the segmented masks and detected corner points, dual-path correspondence matching leverages them to achieve dense and reliable 3D-2D correspondences. Inspired by human visual system, DPCM consists of two pathways, one for correspondence matching of LIP and RGB images, and the other for LDP and depth images. For each pathway, DPCM adopted the approach outlined in [5] to obtain mask matching result from a cost matrix . Each matched mask pair can estimate an affine transformation to guide the correspondence matching. Specifically, we update the corner points in the virtual image to a location that is close to its true projection coordinate using this affine transformation, as follows:
@@
denotes the matching cost between the -th corner point of a mask in the LiDAR virtual image and the -th corner point of a mask in the camera image. (6) consists of structural and textural consistency. The structural consistency measures the structural difference of corner points in the virtual and real image, where denotes the average perimeter of the matched masks and represents the similarity of the neighboring vertices between current and target corner point. The textural consistency derives the relative textural similarity of the neighboring zone. After establishing the cost matrix, a strict criterion is applied to achieve reliable matching. Matches with the lowest costs in both horizontal and vertical directions of are determined as the optimum corner point matches. Since every can trace back to a LiDAR 3D point , and every is related to a pixel (represented in homogeneous coordinates as ) in the camera image, the final correspondence matching result of DPCM is .
III-C Spatial-Temporal Relative Pose Optimization
EdO-LCEC treats the sensor’s operational environment as a spatial-temporal flow composed of multiple scenes across different time instances. In situations with incomplete or sparse point clouds, single-view methods such as [14, 5] are constrained by the limited number of high-quality correspondences. Our environment-driven approach addresses this limitation by integrating multi-view and multi-scene optimization. By merging the optimal matching results from multiple time instances and scenes, this spatial-temporal optimization enhances the selection of correspondences, maximizing the number of high-quality matches and ultimately improving overall calibration accuracy.
Multi-View Optimization. In multi-view optimization, the extrinsic matrix of the -th scene can be formulated as follows:

### v2

@@ Abstract
LiDAR-camera extrinsic calibration (LCEC) is crucial for multi-modal data fusion in mechatronics. Existing methods, whether target-based or target-free, typically rely on customized calibration targets or fixed scene types, limiting their practicality in real-world applications. To address these challenges, we introduce EdO-LCEC, the first environment-driven online calibration approach. Unlike traditional target-free methods, EdO-LCEC observes the feature density of the application environment through a generalizable scene discriminator. Based on this feature density, EdO-LCEC extracts LiDAR intensity and depth features from varying perspectives to achieve higher calibration accuracy. To overcome the challenges of cross-modal feature matching between LiDAR and camera, we propose dual-path correspondence matching (DPCM), which leverages both structural and textural consistency for reliable 3D-2D correspondences. Additionally, our approach models the calibration process as a joint optimization problem utilizing global constraints from multiple views and scenes to enhance accuracy. Extensive experiments on real-world datasets demonstrate that EdO-LCEC outperforms state-of-the-art methods, particularly in sparse or partially overlapping sensor views.
Index Terms:
Environment-driven, LiDAR-camera extrinsic calibration, multi-modal data fusion.
@@ I Introduction
Perception is a critical challenge for autonomous mobile robots [1]. As the front-end module of mechatronic systems, it provides essential environmental information for high-level navigation. Modern LiDAR-camera fusion systems significantly enhance robotic perception capabilities [2, 3, 4, 5]. While LiDARs provide accurate spatial information, cameras capture rich textural details [6]. Their complementary fusion enables robust performance in key robotic tasks including SLAM [7, 8] and object recognition [9]. LiDAR-camera extrinsic calibration (LCEC), which estimates the relative pose between the two sensors, becomes a core and foundational process for effective data fusion in mechatronics.
@@
Current LiDAR-camera extrinsic calibration methods are primarily categorized as either target-based or target-free. Target-based approaches [10, 11, 12, 13, 14, 15] have long been the preferred choice in this field. They are typically offline, relying on customized calibration targets (typically checkerboards). However, they often demonstrate poor adaptability to real-world environments. This is largely because extrinsic parameters may change significantly due to moderate shocks or during extended operations in environments with vibrations. Online, target-free approaches aim at overcoming this problem by extracting informative visual features directly from the environment. Previous works [16, 17] estimate the extrinsic parameters by matching the cross-modal edges between LiDAR projections and camera images. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed edge features. Recent advances in deep learning techniques have spurred extensive explorations, such as [18, 19, 20], to leverage semantic information to aid cross-modal feature matching. Although these approaches have shown compelling performance in specific scenarios, such as urban freeway, they predominantly rely on curated, pre-defined objects, e.g., vehicles, lanes, and traffic poles. On the other hand, several end-to-end deep learning networks [21, 22, 23, 24, 25, 26] have been developed to find a more direct solution for LCEC. While these methods have demonstrated effectiveness on their training datasets like KITTI [27], they highly rely on the training setup and are thus less generalizable. In our previous work, MIAS-LCEC [6], we developed a two-stage coarse-to-fine matching approach using the large vision model MobileSAM [28] to improve cross-modal feature matching. While this method achieves high accuracy with solid-state LiDARs (producing dense point clouds) when the sensors have overlapping fields of view, its performance significantly degrades in challenging scenarios with sparse or incomplete point clouds.
@@
Although target-free LCEC methods eliminate the reliance on calibration targets, they struggle with complex, unstructured scenes due to their inability to adapt to environmental changes. To address this, we introduce environmental observation into target-free calibration, proposing the first environment-driven framework, EdO-LCEC. EdO-LCEC adapts to external conditions to maintain optimal performance by balancing the feature density between LiDAR projection and camera image. In this way, the system is no longer constrained by fixed strategies but can intelligently respond to environmental changes. Specifically, as illustrated in Fig. Environment-Driven Online LiDAR-Camera Extrinsic Calibration, we consider the working environment of the sensors as a sequence composed of multiple scenes. By selectively merging multiple scenes across different times, this approach could achieve high-precision calibration dynamically. A prerequisite for the success of EdO-LCEC is the design of a generalizable scene discriminator. The scene discriminator employs large vision models to conduct depth estimation and image segmentation. In detail, it calculates the feature density of the calibration scene and uses it to guide the generation of multiple virtual cameras for projecting LiDAR intensities and depth. This improved LiDAR point cloud projecting strategy increases the available environmental features, and thus overcomes the previous reliance of algorithms [16, 29] on uniformly distributed geometric or textural features. At each scene, we perform dual-path correspondence matching (DPCM). Different from the C3M in MIAS-LCEC [6], DPCM divides correspondence matching into spatial and textural pathways to fully leverage both geometric and semantic information. Each pathway constructs a cost matrix based on structural and textural consistency, guided by accurate semantic priors, to yield reliable 3D-2D correspondences. Finally, the correspondences obtained from multiple views and scenes are used as inputs for our proposed multi-view and multi-scene joint optimization, which derives and refines the extrinsic matrix between LiDAR and camera. Through extensive experiments conducted on three real-world datasets, EdO-LCEC demonstrates superior robustness and accuracy compared to other SoTA approaches.
@@
EdO-LCEC, the first environment-driven, online LCEC framework that introduces environmental observation into target-free calibration.
Generalizable scene discriminator, which can automatically observe the calibration scene by evaluating the feature density through potential spatial, textural and semantic features extracted by SoTA LVMs.
@@
Multi-view and multi-scene joint relative pose optimization, enabling high-quality extrinsic estimation by integrating multiple perspective views within a single scene and merging distinct scenes across different timestamps.
The remainder of this article is structured as follows: Sect. II reviews SoTA approaches in LCEC. Sect. III introduces EdO-LCEC, our proposed online, target-free LCEC algorithm. Sect. IV presents experimental results and compares our method with SoTA methods. Finally, in Sect. V, we conclude this article and discuss potential future research directions.
@@ II Related Work
Target-based LCEC methods achieve high accuracy using customized calibration targets (typically checkerboards). However, they require offline execution and are significantly limited in dynamic or unstructured environments where such targets are unavailable [30, 20]. Recent studies have shifted to online, target-free approaches to overcome these limitations. Pioneering works [31, 32, 16, 17, 33] estimate the relative pose between the two sensors by aligning the cross-modal edges or mutual information (MI) extracted from LiDAR projections and camera RGB images. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed edges and rich texture, which largely compromise calibration robustness. To circumvent the challenges associated with cross-modal feature matching, several studies [34, 35, 36, 30] have explored motion-based methods. These approaches match sensor motion trajectories from visual and LiDAR odometry to derive extrinsic parameters through optimization. While they effectively accommodate heterogeneous sensors without requiring overlap, they demand precise synchronized LiDAR point clouds and camera images to accurately estimate per-sensor motion, which limits their applicability in real-world scenarios.
Advances in deep learning techniques have driven significant exploration into enhancing traditional target-free algorithms. Some studies [37, 38, 18, 19, 20, 39, 29, 6] explore attaching deep learning modules to their calibration framework as useful tools to enhance calibration efficiency. For instance, [38] accomplishes LiDAR and camera registration by aligning road lanes and poles detected by semantic segmentation. Similarly, [19] employs stop signs as calibration primitives and refines results over time using a Kalman filter. A recent study [29] introduced Direct Visual LiDAR Calibration (DVL), a novel point-based method that utilizes SuperGlue [40] to establish direct 3D-2D correspondences between LiDAR and camera data. On the other hand, several learning-based algorithms [21, 22, 23, 24, 25, 26] try to convert the calibration process into more direct approaches by exploiting end-to-end deep learning networks. Although these methods have demonstrated effectiveness on public datasets like KITTI [27], which primarily focuses on urban driving scenarios, their performance has not been extensively validated on other types of real-world datasets that contain more challenging scenes. Furthermore, their dependence on pre-defined sensor configurations (both LiDAR and camera) poses implementation challenges. In comparison, our approach introduces feature density evaluation to achieve environmental observation. This environment-driven strategy enables robots to maintain high-precision calibration anytime and anywhere.
@@ III Methodology
We first utilize a scene discriminator to perceive the environment through image segmentation and depth estimation, generating virtual cameras that project LiDAR intensity and depth from multiple viewpoints. LiDAR projections from multiple views are further segmented into masks with corner points (Sect. III-A).
The segmented masks with detected corner points are processed along two pathways (spatial and textural) of the dual-path correspondence matching module to establish reliable 3D-2D correspondences (Sect. III-B).
The obtained correspondences are used as inputs for our proposed multi-view and multi-scene joint optimization method, which derives and refines the extrinsic matrix (Sect. III-C).
@@ III-A Generalizable Scene Discriminator
Our environment-driven approach first employs a generalizable scene discriminator to observe the surroundings by generating virtual cameras to project LiDAR point cloud intensities and depth. This discriminator configures both an intensity and a depth virtual camera from the LiDAR’s perspective. This setup yields a LiDAR intensity projection (LIP) image ( and represent the image height and width) and a LiDAR depth projection (LDP) image . To align with the LDP image, the input camera RGB image is processed using Depth Anything V2 [41] to obtain estimated depth images 111In this article, the symbols in the superscript denote the type of target camera ( denotes virtual camera and indicates real camera), and the subscript denotes the source of the image ( is depth and is intensity).. To take advantage of semantic information, we utilize MobileSAM [42] as the image segmentation backbone. The series of detected masks in an image is defined as . The corner points along the contours of masks detected are represented by , where is the total corner points number in the -th mask. An instance (bounding box), utilized to precisely fit around each mask, is centrally positioned at and has a dimension of pixels. To fully exploit the textural information, we evaluate the consistency of corner points using a texture matrix , which encodes intensity values within a local neighborhood. As depicted in Fig. 1, the neighboring vertices of a corner point is defined as . These neighboring vertices are used to calculate structural consistency in dual-path correspondence matching.
@@
where denotes the textural density and represents the structural density. The occlusion challenges caused by the different perspectives of LiDAR and the camera [16, 6], combined with limited feature availability in sparse point clouds, mean that a single pair of virtual cameras is insufficient for a comprehensive view of the calibration scene. Let represent the event of capturing enough effective features, with probability . If we have intensity virtual cameras and depth virtual cameras, the probability of capturing enough effective features is . In theory, as , the probability , leading to . Increasing the number of virtual cameras raises the likelihood of capturing more potential correspondences, thus enhancing calibration accuracy. However, applying an infinite number of virtual cameras during calibration is impractical. Considering the trade-off between calibration accuracy and computational resources, we set the number of multiple virtual cameras to satisfy the balance of feature density:
As depicted in Fig. 2, in practical applications, we set multiple virtual cameras inside a sphere originating from the LiDAR perspective center. Since the feature density is similar if the perspective viewpoints of virtual cameras are close to the initial position, we can assume that and . So and can be obtained as follows:
@@ III-B Dual-Path Correspondence Matching
Given the segmented masks with detected corner points, dual-path correspondence matching leverages them to achieve dense and reliable 3D-2D correspondences. DPCM consists of two pathways, one for correspondence matching of LIP and RGB images, and the other for LDP and depth images. For each pathway, DPCM adopted the approach outlined in [6] to obtain mask matching result from a cost matrix . Each matched mask pair can estimate an affine transformation to guide the correspondence matching. Specifically, we update the corner points in the virtual image to a location that is close to its true projection coordinate using this affine transformation, as follows:
@@
Require:
Textural pathway: Segmented Masks (including detected corner points) obtained from the LIP and RGB image.
Spatial pathway: Segmented Masks (including detected corner points) obtained from the LDP and depth images.
Stage 1 (Reliable mask matching):
(2) Estimate and using (9)-(12) in [6].
Stage 2 (Dense correspondence matching):
(1) For each pathway, update all masks in the virtual image using and .
(2) Construct the corner point cost matrix using (6).
(4) Aggregate all corner point correspondences to form the sets .
denotes the matching cost between the -th corner point of a mask in the LiDAR virtual image and the -th corner point of a mask in the camera image. (6) consists of structural and textural consistency. As illustrated in Fig. 3, the structural consistency measures the structural difference of corner points in the virtual and real image, where denotes the average perimeter of the matched masks and represents the similarity of the neighboring vertices between current and target corner point. The textural consistency derives from the relative textural similarity of the neighboring zone. After establishing the cost matrix, a strict criterion is applied to achieve reliable matching. Matches with the lowest costs in both horizontal and vertical directions of are determined as the optimum corner point matches. Since every can trace back to a LiDAR 3D point , and every is related to a pixel (represented in homogeneous coordinates as ) in the camera image, the final correspondence matching result of DPCM is .
Leveraging the generalizable scene discriminator, the multi-perspective views generated by virtual cameras provide rich spatial-textural descriptors that effectively transform the feature matching problem into a robust 3D-2D correspondence establishment process. DPCM achieves efficient computation by verifying 2D structural and textural consistency instead of performing costly 3D point cloud registration. The pseudo-code of our DPCM is presented in Algorithm 1. During corner point matching within a mask, while the C3M in MIAS-LCEC exclusively considers points from the corresponding mask (i.e., the matched mask associated with the current corner point), our DPCM in EdO-LCEC incorporates additional corner points from neighboring masks. This innovative matching strategy offers two key advantages: (1) it partially eliminates the dependency on dense, high-precision mask matching results, and (2) it substantially avoids errors induced by imperfect mask matching, especially under challenging conditions such as sparse LiDAR point clouds or limited LiDAR-camera FoV overlap.
III-C Multi-View and Multi-Scene Joint Optimization
EdO-LCEC models the sensor’s operational environment as a composition of distinct scenes across different timestamps. While single-view methods (e.g., [16, 6]) suffer from limited high-quality correspondences in sparse or incomplete point cloud scenarios, our environment-driven approach overcomes this constraint through integrating observations from multiple views and scenes. After establishing 3D-2D correspondences by DPCM, EdO-LCEC computes the extrinsic matrix between LiDAR and camera through a multi-view and multi-scene joint optimization. By aggregating optimal correspondences across time and space, our method enhances matching robustness, maximizes high-quality feature associations, and ultimately improves calibration accuracy.
In multi-view optimization, the extrinsic matrix of the -th scene can be formulated as follows: